OpenStudy (anonymous):

A line passes through point (3, 7) and has a slope of 34. What is the equation of the line?

OpenStudy (anonymous):

slope of 3/4 **

OpenStudy (anonymous):

you can use point slope form \[(y-y_1) = m (x-x_1)\] where (x1,y1) = given points m = slope

OpenStudy (anonymous):

I have options y=3/4x+19/4 or y=3x-19

OpenStudy (anonymous):

\[y- 7 = \frac{ 3 }{ 4 }*(x -3)\] \[y= \frac{ 3x }{ 4 } - \frac{ 9 }{ 4 } +7\] \[y = \frac{ 3x }{ 4 } - \frac{ 9+ (4*7) }{ 4 }\] \[y = \frac{ 3x }{ 4 } - \frac{ 37 }{ 4 }\]

OpenStudy (anonymous):

\[y = \frac{ 3x }{ 4 } + \frac{- 9+ (4*7) }{ 4 }\] \[y = \frac{ 3x }{ 4 } + \frac{19 }{ 4 }\]

OpenStudy (anonymous):

I don't get it

OpenStudy (anonymous):

I wrote wrong ans equation in calculation part correct answer is y = 3/4(x) +19/4

OpenStudy (anonymous):

If point A(x, 5) lies on the line, the value of x is 33 or 1/3?

OpenStudy (anonymous):

for same equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

given y= 5, we have to find x \[5 = \frac{ 3x }{ 4 } +\frac{ 19 }{ 4 }\] \[5 -\frac{ 19 }{ 4 } = \frac{ 3x }{ 4 } \] \[\frac{(20 - 19) }{ 4 } = \frac{ 3x }{ 4 } \] \[\frac{1 }{ 4 } *\frac{ 4 }{ 3 } = x \] \[\frac{1 }{ 3 } = x \]